OpenStudy (anonymous):

solve f(x) = ax^3+bx^2+cx+d

OpenStudy (anonymous):

is the same, think, same solution as the answer because in f(x). x doesnt some other value. or the equation doesnt have any like terms to simplyfy.... so f(x) = ax^3+bx^2+cx+d will equal f(x) = ax^3+bx^2+cx+d

OpenStudy (anonymous):

x doesnt have * another value other then just x

OpenStudy (anonymous):

I don't think you can solve that unless you make it like: solve f(x) = ax^3+bx^2+cx+d = 0, in which you would use the quadratic formula.